OpenStudy (anonymous):

Part 1: Solve the inequality 5x − 2 2x + 10 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

Is that 5x − 2<= 2x + 10 ?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

Add 2, remove 2x from the equation, and divide everything by 5.

OpenStudy (anonymous):

what bout part 2

OpenStudy (anonymous):

Okay, so once you've solved for x, you have to realize what the sign of the "=" or whatever is saying. = just means it's the equation. However, when you have <=, it means the curve of x is less than or equal to "?" The "equal to" part means you'll shade the curve, and the less than part means you'll shade the region BELOW the curve. If it said >=, you would shade the curve and ABOVE the curve.

OpenStudy (anonymous):

Nanoman is correct...add 2 to both sides, this will cancel the 2 on the left side. Subtract 2x from both sides, this will cancel the 2x on the right side. Then divide by 5 to get x by itself.

OpenStudy (anonymous):

i got x = 2.8

OpenStudy (anonymous):

I got something different. Can you type your work?

OpenStudy (anonymous):

5x - 2 <= 2x + 10 + 2 +2 5x <= 4x + 10 5x <= 14x --------- = x = 2.8 5 5

OpenStudy (anonymous):

no.... 5x - 2 <= 2x + 10 +2 + 2 ----------------- 5x <= 2x + 12 -2x -2x -------------- 3x <= 12 oops....you are supposed to divide by 3 not 5. 3x/3 <= 12/3 x = 4

OpenStudy (anonymous):

You forgot to add the 2 to the 10, not the 2x. That's what messed you up.

OpenStudy (anonymous):

well thank you guys :)